### v1

@@ 6 Search strategy
To maximise the sensitivity and to fully exploit the kinematic differences between the IDM and the SM backgrounds, a Neural Network (NN) is used to separate signal from background after the baseline selections. The kinematic features used as input into the model are similar to those used in Refs. Kalinowski:2018kdn,Zarnecki:2020swm which include both low-level and high-level, derived, event features.
@@
The dilepton pair $E_{\ell\ell}$ and invariant mass $M_{\ell\ell}$ distributions are shown in Figs. 7 and 8 for the different background processes and several representative signal points, for the $\mu\mu$ selection and at $\sqrt{s}=$240 and 365 GeV, respectively. Note that the simulation curve shown for each signal point is the sum of both the $HH\ell\ell$ and $HH\ell\ell\nu\nu$ processes. In these distributions, large variations are seen in kinematic shapes between different signal points. Therefore, to take into account these differences, the network used is a parametric Neural Network Baldi:2016fzo,Anzalone:2022hrt (pNN) in which the two signal masses, $M_{H}$ and $M_{A}$, are also used as input. This allows for training of only a single network whilst ensuring strong sensitivity to all signal points in the parameter space. For background events, which have no intrinsic IDM parameters associated with them, $M_{H}$ and $M_{A}$ are randomly chosen at the start of every training epoch from the set of IDM masses trained on. More details on pNNs and definitions of the relevant vocabulary can be found in Refs. Baldi:2016fzo,Anzalone:2022hrt.
Figure 7: Distribution of the dimuon pair energy $E_{\mu\mu}$ (left) and invariant mass $M_{\mu\mu}$ (right) for $\sqrt{s}=$240 GeV, for the SM backgrounds and several selected signal mass points for the scenario S1. The signal mass points are labeled by IDM $M_{H}$/$M_{A}$ with masses given in GeV.
Figure 8: Distribution of the dimuon pair energy $E_{\mu\mu}$ (left) and invariant mass $M_{\mu\mu}$ (right) for $\sqrt{s}=$365 GeV, for the SM backgrounds and several selected signal mass points for the scenario S1. The signal mass points are labeled by IDM $M_{H}$/$M_{A}$ with masses given in GeV.
The pNN is implemented in PyTorch pytorch, and contains four sequential feed-forward layers each with 250 units and a ReLU activation function, totaling roughly 200,000 learnable parameters. The last three layers are preceded by a dropout layer with a rate of $0.2$, and the network output is passed through a sigmoid function for binary classification. The loss function for the training of the model is a weighted binary cross-entropy loss function:
Weighted $\text{Weighted BCE}=-\frac{1}{N}\sum_{i=0}^{N}w_{i}[y_{i}\log(\text{pred}_{i})+(1-y_{i})\log(1-\text{pred}_{i})]$ (3)
@@
The pNN can smoothly interpolate to mass points not trained on when applied to the background samples. The signal distributions at intermediate points are found using cubic spline interpolation. This involves a shape parameter $\varepsilon$ that is determined by minimising the leave-one-out cross-validation error Rippa1999AnAF. The interpolation capability of the setup was checked for several points of the parameter space, by comparing the results when those points are included or not in the training of the model. The pNN results were also compared with a simple implementation of a boosted decision tree using the same input features inside the TMVA framework TMVA, and comparable results were obtained on the individual IDM BPs.
The pNN distribution is shown for several representative signal mass points in Fig. 9 and 10 for $\sqrt{s}=$240 and 365 GeV, respectively. The top row shows the full distribution whilst the bottom row shows the distribution above 0.9 which is used for extracting the results in the next section.
Figure 9: Output discriminator of the pNN for several signal points for $\sqrt{s}=$240 GeV samples summing $ee$ and $\mu\mu$ contributions if $M_{A}-M_{H}$$>30$GeV, only $\mu\mu$ elsewhere. Top: full-range distribution, bottom: selection defining the signal region for the extraction of the results.
Figure 10: Output discriminator of the pNN for several signal points for $\sqrt{s}=$365 GeV samples summing $ee$ and $\mu\mu$ contributions if $M_{A}-M_{H}$$>30$GeV, only $\mu\mu$ elsewhere. Top: full-range distribution, bottom: selection defining the signal region for the extraction of the results.
@@ 7 Results
A maximum likelihood fit of the pNN output above 0.9 is performed using the CMS Combine package CMS:2024onh. A 95% confidence-level (CL) upper limit is derived using the asymptotic approximation Cowan:2010js. A single source of uncertainties is included as nuisance parameters: the MC statistical uncertainties Conway:2011in. Future work should also include systematic uncertainties as additional nuisance parameters.
@@
The $ee$ and $\mu\mu$ channels are fitted simultaneously. As the inclusive $ee$ production was simulated only for $M_{ee}>30$ GeV, the two final states $ee$ and $\mu\mu$ are combined for $M_{A}-M_{H}$$>30$ GeV, below which only the $\mu\mu$ result is shown.
@@
The 95% CL exclusion region obtained is shown in Fig. 11 (left) in the $M_{A}-M_{H}$ vs $M_{H}$ plane, for $\sqrt{s}=$240 GeV, for the total integrated luminosity of 10.8 ab-1. The three different scenarios defined in section 2 are shown as separate lines, well within the ±$\pm 1\sigma$ contour shown in green for scenario S1, including only the statistical uncertainties. The kinematic limit for on-shell production is highlighted by the dashed line. With the full luminosity of 10.8 ab-1, almost the entire parameter space can be excluded at 95% CL, reaching $M_{H}$$=110$GeV.
@@
Figure 11: 95% CL expected exclusion (left) and $>5$-sigma discovery region (right) in the $M_{A}-M_{H}$ vs $M_{H}$ plane, for $\sqrt{s}=$240 GeV with total integrated luminosities of 10.8 and 2 ab-1.
The $>5$-sigma discovery region is shown in Fig. 11 (right) for $\sqrt{s}=$240 GeV, in the scenario S1, for the total integrated luminosity of 10.8 ab-1 and 2 ab-1. The regions excluded by relic density constraints and LEP SUSY recast (see section 2) are overlayed. The discovery reach extends to $M_{H}$$=108\,\text{G\text{eV}}$ for $M_{A}-M_{H}$$=15$GeV, for the nominal FCC luminosity scenario.
@@
The 95% CL limit (left) and $>5$-sigma discovery (right) regions are shown in Fig. 12 for $\sqrt{s}=$365 GeV, for the total integrated luminosity of 2.7 ab-1. The regions excluded by relic density constraints and LEP SUSY recast are overlayed. As for $\sqrt{s}=$240 GeV, the three different scenarios defined in section 2 are shown as separate lines and give results within the ±$\pm 1\sigma$ contour shown in green for scenario S1, including only the statistical uncertainties. The wedge in the distribution observed for a mass splitting around the on-shell Z boson mass is due to reduced sensitivity from reduced discrimination against background. The 95% CL excluded region reaches $M_{H}$$=165$GeV. The discovery reach extends to $M_{H}$$=157$GeV for $M_{A}-M_{H}$$=15$GeV for the nominal FCC luminosity scenario, and is also shown in Fig. 12 right for a scenario with ten times less luminosity.
@@
Figure 12: 95%CL limit (left) and $>5$-sigma discovery contour (right) in the $M_{A}-M_{H}$ vs $M_{H}$ plane, for $\sqrt{s}=$365 GeV with a total integrated luminosity of 2.7 ab-1.
@@ 8 Conclusion
The Inert Doublet Model predicts additional scalars which couple only to bosons. The lighest neutral scalar, $H$, is also stable and provides an adequate dark matter candidate. The pair production of such new scalars is investigated in a final state containing two electrons or two muons, with $e^{+}e^{-}$ collisions at $\sqrt{s}=$240 and 365 GeV, in the context of the future circular collider at CERN, FCC-ee. With a total integrated luminosity of 10.8 (2.7)  ab-1for $\sqrt{s}=$240 (365) GeV, almost the entire parameter space available in the $M_{A}-M_{H}$ vs $M_{H}$ plane is expected to be excluded at 95% CL, reaching up to $M_{H}$$=110\,(165)$GeV, after using a parametric neural network trained to discriminate the different signal mass points against the backgrounds. The discovery reach is also explored and covers again almost the entire kinematic phase space available at $\sqrt{s}=$240 (365) GeV reaching $M_{H}$$=108\,(157)\,\text{G\text{eV}}$ for $M_{A}-M_{H}$$=15$GeV.
Acknowledgements
TR has received support from grant number HRZZ-IP-2022-10-2520 from the Croatian Science Foundation (HRZZ); in addition, this work was supported by the Science and Technology Facilities Council, UK. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). The authors thank A.F. Zarnecki for useful discussions. TR also wants to thank the Kavli Institute for Theoretical Physics (KITP) at the University of California, Santa Barbara, and in particular the organizers of the program "What is particle theory ?", for their hospitality while parts of this work were completed.

### v2

@@ 6 Search strategy
To maximise the sensitivity and to fully exploit the kinematic differences between the IDM and the SM backgrounds, a Neural Network (NN) is used to separate signal from background after the baseline selections. The kinematic features used as input into the model are similar to those used in Refs. [12, 13] which include both low-level and high-level, derived, event features.
@@
The dilepton pair $E_{\ell\ell}$ and invariant mass $M_{\ell\ell}$ distributions are shown in Figs. 7 and 8 for the different background processes and several representative signal points, for the $\mu\mu$ selection and at $\sqrt{s}=$240 and 365 GeV, respectively. Note that the simulation curve shown for each signal point is the sum of both the $HH\ell\ell$ and $HH\ell\ell\nu\nu$ processes. In these distributions, large variations are seen in kinematic shapes between different signal points. Therefore, to take into account these differences, the network used is a parametric Neural Network [46, 47] (pNN) in which the two signal masses, $M_{H}$ and $M_{A}$, are also used as input along with the kinematic features listed above. By including the signal parameters as input, the model learns the optimal selection as a function of which mass hypothesis is being tested. As the signals can have different kinematic distributions, the trained pNN will apply a different set of selections that are optimal for each mass point. This allows for training of only a single network per centre-of-mass energy whilst ensuring strong sensitivity to all signal points in the parameter space. For background events, which have no intrinsic IDM parameters associated with them, $M_{H}$ and $M_{A}$ are randomly chosen at the start of every training epoch from the set of IDM masses trained on. More details on pNNs and definitions of the relevant vocabulary can be found in Refs. [46, 47].
Figure 7: Distribution of the dimuon pair energy $E_{\mu\mu}$ (left) and invariant mass $M_{\mu\mu}$ (right) for $\sqrt{s}=$240 GeV, for the SM backgrounds and several selected signal mass points for the scenario S1. The signal mass points are labeled by IDM $M_{H}$$M_{A}$ with masses given in GeV.
Figure 8: Distribution of the dimuon pair energy $E_{\mu\mu}$ (left) and invariant mass $M_{\mu\mu}$ (right) for $\sqrt{s}=$365 GeV, for the SM backgrounds and several selected signal mass points for the scenario S1. The signal mass points are labeled by IDM $M_{H}$$M_{A}$ with masses given in GeV.
The pNN is implemented in PyTorch [48], and contains four sequential feed-forward layers each with 250 units and a ReLU activation function, totaling roughly 200,000 learnable parameters. The last three layers are preceded by a dropout layer with a rate of $0.2$, and the network output is passed through a sigmoid function for binary classification. The loss function for the training of the model is a weighted binary cross-entropy loss function:
$\text{Weighted BCE}=$
$-\frac{1}{N}\sum_{i=0}^{N}w_{i}[y_{i}\log(\text{pred}_{i})+(1-y_{i})\log(1-\text{pred}_{i})]$
@@
The pNN can smoothly interpolate to mass points not trained on when applied to the background samples. The signal distributions at intermediate points are found using cubic spline interpolation. This involves a shape parameter $\varepsilon$ that is determined by minimising the leave-one-out cross-validation error [49]. The interpolation capability of the setup was checked for several points of the parameter space, by comparing the results when those points are included or not in the training of the model. The pNN results were also compared with a simple implementation of a boosted decision tree using the same input features inside the TMVA framework [50], and comparable results were obtained on the individual IDM BPs.
The pNN distribution is shown for several representative signal mass points in Fig. 9 and 10 for $\sqrt{s}=$240 and 365 GeV, respectively. The top row shows the full distribution whilst the bottom row shows the distribution above 0.9 which is used for extracting the results in the next section. The efficiency of the selection pNN >> 0.9 is above 50% for all signal points, and above 90% in the region $10\,<\,M_{A}-M_{H}\,<\,50\,\text{G\text{eV}}$. The lowest efficiencies are obtained when $M_{A}-M_{H}$ is close to the $Z$ boson mass, and at very low $M_{A}-M_{H}$.
Figure 9: Output discriminator of the pNN for several signal points for $\sqrt{s}=$240 GeV samples summing eeee and $\mu\mu$ contributions if $M_{A}-M_{H}$>30>30GeV, only $\mu\mu$ elsewhere. Top: full-range distribution, bottom: selection defining the signal region for the extraction of the results.
Figure 10: Output discriminator of the pNN for several signal points for $\sqrt{s}=$365 GeV samples summing eeee and $\mu\mu$ contributions if $M_{A}-M_{H}$>30>30GeV, only $\mu\mu$ elsewhere. Top: full-range distribution, bottom: selection defining the signal region for the extraction of the results.
@@ 7 Results
A maximum likelihood fit of the pNN output above 0.9 is performed using the CMS Combine package [51]. The threshold of 0.9 is chosen to have maximum signal significance over backgrounds for all points. Both 95% confidence level (CL) expected upper limits, assuming no signal for exclusion, and expected significance, assuming a signal rate predicted at a given parameter point for discovery, are calculated using profiled likelihood ratio test statistics, using asymptotic approximations for their distributions [52]. A single source of uncertainties is included as nuisance parameters: the MC statistical uncertainties [53]. Future work should also include systematic uncertainties as additional nuisance parameters.
@@
The $ee$ and $\mu\mu$ channels are fitted simultaneously. As the inclusive $ee$ production was simulated only for $M_{ee}>30$ GeV, the two final states $ee$ and $\mu\mu$ are combined for $M_{A}-M_{H}$>30>30 GeV, below which only the $\mu\mu$ result is shown.
@@
Contours are drawn for when the expected upper limit is smaller than the theoretical cross-section, giving the 95% CL exclusion region shown in Fig. 11 (left) in the $M_{A}-M_{H}$ vs $M_{H}$ plane, for $\sqrt{s}=$240 GeV, for the total integrated luminosity of 10.8 ab-1. The three different scenarios defined in section 2 are shown as separate lines, well within the $\pm 1\sigma$ contour shown in green for scenario S1, including only the statistical uncertainties. The kinematic limit for on-shell production is highlighted by the dashed line. With the full luminosity of 10.8 ab-1, almost the entire parameter space can be excluded at 95% CL, reaching $M_{H}$=110=110GeV.
@@
Figure 11: 95% CL exclusion contour (left) and >5>5-sigma discovery contour (right) in the $M_{A}-M_{H}$ vs $M_{H}$ plane, for $\sqrt{s}=$240 GeV with total integrated luminosities of 10.8 and 2 ab-1. Only statistical uncertainties have been included in the fit.
Contours are also drawn for when the expected significance reaches $5$-sigma, giving the >5>5-sigma discovery region shown in Fig. 11 (right) for $\sqrt{s}=$240 GeV, in the scenario S1, for the total integrated luminosity of 10.8 ab-1 and 2 ab-1. The regions excluded by relic density constraints and LEP SUSY recast (see section 2) are overlayed [^5] Note that in principle scenarios 2 and 3 are now excluded by direct detection constraints from [29]. However, the main target here was to check the impact of switching on the respective channel with a 125 GeV  scalar mediator. We therefore chose to keep the corresponding results for illustration purposes.. The discovery reach extends to $M_{H}$$108\,\text{G\text{eV}}$ for $M_{A}-M_{H}$=15=15GeV, for the nominal FCC luminosity scenario. The other luminosity scenario of 2 ab-1 is considered for comparison with a linear collider option at $\sqrt{s}=$250 GeV [54].
@@
The 95% CL exclusion (left) and >5>5-sigma discovery (right) regions are shown in Fig. 12 for $\sqrt{s}=$365 GeV, for the total integrated luminosity of 2.7 ab-1. The regions excluded by relic density constraints and LEP SUSY recast are overlayed. As for $\sqrt{s}=$240 GeV, the three different scenarios defined in section 2 are shown as separate lines and give results within the $\pm 1\sigma$ contour shown in green for scenario S1, including only the statistical uncertainties. The wedge in the distribution observed for a mass splitting around the on-shell Z boson mass is due to reduced sensitivity from reduced discrimination against background. The 95% CL excluded region reaches $M_{H}$=165=165GeV. The discovery reach extends to $M_{H}$=157=157GeV for $M_{A}-M_{H}$=15=15GeV for the nominal FCC luminosity scenario, and is also shown in Fig. 12 right for a scenario with ten times less luminosity , corresponding to a linear collider option at $\sqrt{s}=$350 GeV [54].
@@
Figure 12: 95%CL exclusion contour (left) and >5>5-sigma discovery contour (right) in the $M_{A}-M_{H}$ vs $M_{H}$ plane, for $\sqrt{s}=$365 GeV with a total integrated luminosity of 2.7 ab-1. Only statistical uncertainties have been included in the fit.
@@ 8 Conclusion
The Inert Doublet Model predicts additional scalars which couple only to bosons. The lighest neutral scalar, $H$, is also stable and provides an adequate dark matter candidate. The pair production of such new scalars is investigated in a final state containing two electrons or two muons, with $e^{+}e^{-}$ collisions at $\sqrt{s}=$240 and 365 GeV, in the context of the future circular collider at CERN, FCC-ee. A parametric neural network is trained to discriminate the different signal mass points against the backgrounds. A maximum likelihood fit of the output of the neural network is performed, including only statistical uncertainties. With a total integrated luminosity of 10.8 (2.7)  ab-1 for $\sqrt{s}=$240 (365) GeV, the discovery reach is explored and covers almost the entire kinematic phase space available at $\sqrt{s}=$240 (365) GeV reaching $M_{H}$$108\,(157)\,\text{G\text{eV}}$ for $M_{A}-M_{H}$=15=15GeV. Furthermore, almost the entire parameter space available in the $M_{A}-M_{H}$ vs. $M_{H}$ plane is expected to be excluded at 95% CL, reaching up to $M_{H}$=110(165)=110\,(165)GeV. Parts of the 2-dimensional plane are already excluded by a recast of LEP results and constraints from relic density, i.e. overclosure of the universe. The FCC-ee reach will be complementary to that of higher-energy collisions from the LHC and HL-LHC, which probe an area with larger $M_{A}-M_{H}$.
[…]
Acknowledgments
TR has received support from grant number HRZZ-IP-2022-10-2520 from the Croatian Science Foundation (HRZZ); in addition, this work was supported by the Science and Technology Facilities Council, UK. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). The authors thank A.F. Zarnecki for useful discussions. TR also wants to thank the Kavli Institute for Theoretical Physics (KITP) at the University of California, Santa Barbara, and in particular the organisers of the programme ”What is particle theory ?”, for their hospitality while parts of this work were completed.
Declarations
  • Competing interests: the authors have no relevant financial or non-financial interests to disclose.
  • Data availability: this manuscript has no data associated beyond references given in the text.